OpenStudy (curry):

x-6i=(3-i)+(4-2yi) x-6i=(3-i)+(4-2yi) @Mathematics

OpenStudy (anonymous):

Could you give us some context? What are we solving for?

OpenStudy (anonymous):

Is i \[\sqrt{-1}\] or a variable?

OpenStudy (curry):

i = square root of -1

OpenStudy (anonymous):

And what are we solving for?

OpenStudy (curry):

x and y

OpenStudy (phi):

the trick to these is to realize that you can treat the real and the imaginary parts as two separate problems. For example, the real parts give you this equation: x= 3+4

OpenStudy (anonymous):

X- 6i = 7-3yi compring tow equations X= 7 -3y = -6 y= 2